OpenStudy (anonymous):

Chris is playing a game with his father with a six-sided die. What is the expected value per roll that Chris will get if his father agrees to give him $2 multiplied by every even number rolled and $1 for every odd number rolled? $1.50 $4.50 $10.50 $11.00 $12.00

OpenStudy (anonymous):

its (1/6)(2+2+2+1+1+1) =9/6 =$1.50